OpenStudy (anonymous):

680cm= ? ft give answer in decimal form to nearest tenth

OpenStudy (anonymous):

1cm = 0.03ft So multiply 0.03 by 680 and voila.

OpenStudy (anonymous):

30.48cm = 1ft Divide 680 by 30.48

OpenStudy (kropot72):

2.54 cm = 1 inch 2.54 * 12 cm = 1 ft \[\frac{680}{2.54\times 12} = ? ft\]